Three gases (8.00 g of methane, C H 4 , 18.0 g of ethane, C 2 H 6 , and an unknown amount of propane, C 3 H 8 ) were added to the same 10.0- L container. At 23.0 ∘ C , the total pressure in the container is 5.30 atm . Calculate the partial pressure of each gas in the container. Express the pressure values numerically in atmospheres, separated by commas. Enter the partial pressure of methane first, then ethane, then propane.

A gaseous mixture of O2 and N2 contains 30.8 % nitrogen by mass. What is the partial pressure of oxygen in the mixture if the total pressure is 605 mmHg ?

Express you answer numerically in millimeters of mercury.
Answer #1

Solution:

According to ideal gas equation,

PV = n RT

n = PV/RT

Where, P = 5.30 atm

V = 10 L

R = 0.0821 L atm K-1 mol-1

T = 23 + 273 = 296 K

Thus,

n = 5.30 atm x 10 L / 0.0821 atm L K-1 mol-1 x 296 K

n = 2.18 mol

Total number of moles = 2.18 mol

Number of moles of methane = mass / molar mass

= 8 g / 16 g mol-1 = 0.50 mol

Number of moles of ethane = mass / molar mass

= 18 g / 30 g mol-1 = 0.60 mol

Thus, number of moles of C3H8= 2.18 mol - (0.50 + 0.60)

= 2.18 - 1.10 = 1.08 mol

According to Raults, law:

Partial pressure = mole fraction x Total pressure

= (number of moles / Total number of moles) x Total pressure

Thus,

For methane:

pCH4 = (0.50 / 2.18) x 5.30 atm = 1.22 atm

For ethane:

pC2H6 = (0.60 / 2.18) x 5.30 atm = 1.46 atm

For propane:

pC3H8 = (1.08 / 2.18) x 5.30 atm = 2.62 atm